OpenStudy (anonymous):

How to write (2n)! in general form?

OpenStudy (anonymous):

"general form"?

OpenStudy (anonymous):

i'm guessing expaneded form?

OpenStudy (anonymous):

yeap

OpenStudy (vijay):

2 [ n * (n-1) -------- 3, 2 1 ]

OpenStudy (anonymous):

But I want to include n! here in the expanded form

OpenStudy (anonymous):

\[(2n)!=2n(2n-1)(2n-2)...\times n!\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=(2n)!

OpenStudy (anonymous):

Why is it satellite? Though I know that n!=n*(n-1)*........2*1

OpenStudy (anonymous):

I want to know how?

OpenStudy (anonymous):

What about [2(n+1)]! then

OpenStudy (anonymous):

(2n)! isnt 2n! (2n)! is (2n)(2n-1)(2n-2)...(n+1)(n)(n-1)...(2)(1) Examples: n = 3 (2n)! = (2*3)! = 6! = 720 2*3! = 2*6 = 12 totally different.